# Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark

### v2

@@ A.1 UCR Time Series Classification datasets
A.2 Proposed, non z-normalized, datasets
Train Test Length Class Type
Data
[…]
Chinatown 20 345 24 2 Traffic
Crop 7200 16800 46 24 Image
DodgerLoopDay 78 80 288 7 Sensor
EOGVerticalSignal 362 362 1250 12 EOG
GestureMidAirD1 208 130 360 26 Trajectory
GunPointAgeSpan 135 316 150 2 Motion
HouseTwenty 34 101 3000 2 Device
InsectEPGRegularTrain 62 249 601 3 EPG
MelbournePedestrian 1200 2450 24 10 Traffic
PLAID 537 537 Vary 11 Device
Rock 20 50 2844 4 Spectrum
SemgHandGenderCh2 300 600 1500 2 Spectrum
SmoothSubspace 150 150 15 3 Simulated
UMD 36 144 150 3 Simulated
[…]
AluminiumConcentration 440 189 2542 2 Environnement
BitcoinSentiment 232 100 24 2 Sentiment
ChilledWaterPredictor 321 138 168 2 Energie
CopperConcentration 440 189 2542 2 Environnement
Covid19Andalusia 142 62 91 2 Santé
DailyOilGasPrices 133 58 30 2 Économie
DhakaHourlyAirQuality 1447 621 24 2 Environnement
ElectricityPredictor 567 243 168 2 Energie
FloodModeling3 429 184 266 2 Environnement
HouseholdPowerConsumption1 1001 430 1440 2 Energie
HotwaterPredictor 245 106 168 2 Energie
MadridPM10Quality_nmv 4845 2077 168 2 Environnement
ParkingBirmingham_eq 1391 597 14 2 Environnement
PrecipitationAndalusia_nmv 470 202 365 2 Environnement
SierraNevadaMountainsSnow 350 150 30 2 Environnement
SolarRadiationAndalusia_nmv 470 202 365 2 Energie
SteamPredictor 210 90 168 2 Energie
TetuanEnergyConsumption 254 110 144 2 Energie
WindTurbinePower 596 256 144 2 Energie
Table 3: New datasets collection : 35 datasets from both the UCR archive (dashed line) and the Monash UEA extrinsic regression archive. When missing values and/or varying lengths, replace missing values with 0 and pad series to maximum length with 0. All of the datasets are not $z$-normalized and thus don’t suffer from potential data leakage. Italic datasets are not included when classes are imbalanced as problems become too difficult for the chosen classifiers.
B Supplementary results
B.1 Additional figures : Standard cost setting
Figure 8: Standard cost setting, non $z$-normalized proposed datasets (orange cylinder). Compared to Figure 3(a), the global ranking is not altered much. One can observe that for $\alpha\in[0.5,0.7]$ the top group is now more populated, gathering the first six approaches, probably due to the limited amount of datasets available in this case.
B.2 Additional figures : Anomaly detection cost setting
Figure 9: Anomaly detection cost setting, original UCR datasets (blue cylinder). Compared to Figure 6(a), one can see that Calimera is now clearly dominating all other methods for all $\alpha$. The global ranking remains globally stable otherwise.
B.3 Removing calibration
Figure 10: Standard cost setting, original UCR datasets (blue cylinder). The calibration step is now removed, i.e. the outputs from the decision function is now simply passed through a softmax function. Both Calimera and Proba Threshold suffer heavily from using uncalibrated scores.
Figure 11: Pairwise comparison (?), calibration ($calib$) vs no calibration ($no\>\>calib$), $\alpha=0.8$. Square colors are indexed on the mean AvgCost difference. For example, Calimera has a lower mean AvgCost when trained over calibrated scores: it appears in dark blue. The Wilcoxon p-value is equal to 0.0288, which is lower than significance level equal to 0.05. Thus, Calimera statistically under-performs when using uncalibrated scores. This is also the case for the Proba Threshold method.
B.4 Changing the base classifier
Figure 12: Standard cost setting, original UCR datasets (blue cylinder). The base classifier is now Weasel 2.0 (?). Results are very close to those exposed in Figure 3(a).
Figure 13: Standard cost setting, original UCR datasets (blue cylinder). The base classifier is now a pipeline including features extraction with tsfresh (?) and classification using XGBoost (?). Results are a bit noisier than those exposed in Figure 3(a).
B.5 Impact of z-normalization
@@
Clearly, using z-normalized datasets is not applicable in practice, as it would require knowledge of the entire incoming time series. In a research context, previous work has used such training sets to test the proposed algorithms. Our goal here, is to assess whether this could have a large impact on the performances. For example, when a normalized time series has a low variance at the beginning, we can expect a high variance in the rest of the series since the mean variance is 1. There is therefore an information leakage that can be exploited by a ECTS algorithm, while this is not representative of what happens in real applications. A proposal such as the one presented in (?), where the z-normalization of available time series is repeated at each time step, has its own problems. In particular, it means that if a single classifier is used for all time steps, the representation of $\mathbf{x}_{t}$ can be different at times $t$ and $t+1$ and all further time steps which can induce confusion for the classifier.
@@
On the one hand, z-normalization induces an information leakage that could help methods to unduly exploit knowledge about the future of incoming time series. On the other hand, any normalization rescale the signal and therefore, potentially, hinder the recognition of telltale features. So, does z-normalization affect the performance of ECTS methods? And if yes, in which way?
@@
In order to answer this question, we took the new datasets collection described in Section 4.1. They are indeed not z-normalized originally. We duplicate and z-normalized them to get a second collection. As explained in Section 4.1, some extrinsic regression datasets has been converted into classification one. Here, the threshold value chosen to discretize the output into binary classes has been set the median of the regression target. In this way, classes within those datasets are equally populated.
@@
In these experiments, the delay cost is linear as in Section 4.2 and as in most of the literature. Figure 14 reports pairwise comparisons done on the 35 datasets. We look at $\alpha=0.8$, as this is the only value for which significant differences are observed. One can see that most of the trigger models do not actually benefit from the z-normalization. Quite the opposite: out of nine trigger models, only one, i.e. Ecdire, actually has a better mean AvgCost when being trained on z-normalized data. Regarding the remaining methods, both Stopping Rule and TeaserAvg performs significantly worse when operating on z-normalized data. Those trends are quite similar for other $\alpha$ values, without any significance on the statistical tests though. Thus, while z-normalization has some impact, since privileged information from the future can be leaked, our experiments, for the proposed datasets collection at least, show that this does not alter the overall results reported in the literature, and are globally in accordance with the results presented in Section 4.
Figure 14: Pairwise comparison (?), $z$-normalization ($Z$) vs no $z$-normalization ($\bar{Z}$), $\alpha=0.8$. Square colors are indexed on the mean AvgCost difference. For example, Calimera has a lower mean AvgCost when trained over non $z$-normalized datasets by 1.3e-2 and appears in light blue. It beats the $z$-normalized version over 25 datasets, loses over 9 and are tied on 1. The Wilcoxon p-value is equal to 0.2224, which is higher than significance level equal to 0.05. Thus, no statistical difference can be observed for the considered approach.
Figure 15: Multi-comparison-matrices (?). The upper triangle, with dark blue contours, displays the comparison of the competitive methods trained over non z-normalized dataset (orange cylinder). The values within this triangle has to be read by lines, i.e. for a considered line, red shades indicate better performances, blue shades weaker performances. The lower triangle, with dark red contours, is the comparison of the methods trained over the same datasets, z-normalized (chocolate cylinder). The values within this triangle has to be read by columns, i.e. for a considered column, red shades indicate better performances, blue shades weaker performances. The complete figure being symmetrical indicates that z-normalization does not impact much relative ranking between methods.
B.6 Anomaly detection cost setting : an ablation study
@@
Figure 16: Exponential delay cost, symmetric binary misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).
@@
Figure 17: Linear delay cost, non symmetric imbalanced misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).
@@
Figure 18: Linear delay cost, symmetric binary misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).

### v6

@@ E.1 UCR Time Series Classification datasets
E.2 Proposed, non z-normalized, datasets
Size Length Class Type AUC Gain AUC Gain
train test
Data (half/full) (half/full)
BME 180 128 3 Simulated (7%/7%) (11%/16%)
Chinatown 365 24 2 Traffic (1%/1%) (0%/0%)
Crop 24000 46 24 Image (9%/10%) (8%/9%)
DodgerLoopDay 158 288 7 Sensor (4%/5%) (14%/17%)
EOGVerticalSignal 724 1250 12 EOG (35%/35%) (43%/45%)
GestureMidAirD1 338 360 26 Trajectory (11%/12%) (23%/26%)
GunPointAgeSpan 451 150 2 Motion (7%/7%) (7%/7%)
HouseTwenty 135 3000 2 Device (1%/1%) (5%/6%)
MelbournePedestrian 3650 24 10 Traffic (15%/15%) (15%/16%)
PLAID 1074 Vary 11 Device (0%/0%) (2%/2%)
Rock 70 2844 4 Spectrum (1%/1%) (12%/12%)
SemgHandGenderCh2 900 1500 2 Spectrum (1%/2%) (11%/13%)
SmoothSubspace 300 15 3 Simulated (21%/37%) (20%/38%)
UMD 180 150 3 Simulated (10%/11%) (22%/28%)
AcousticContaminationMadrid 138 365 2 Environment (1%/2%) (5%/7%)
AluminiumConcentration 629 2542 2 Environment (3%/4%) (5%/10%)
BitcoinSentiment 332 24 2 Sentiment (11%/13%) (-5%/-6%)
ChilledWaterPredictor 459 168 2 Energy (0%/0%) (8%/11%)
CopperConcentration 629 2542 2 Environment (4%/5%) (4%/3%)
Covid19Andalusia 204 91 2 Health (7%/8%) (7%/17%)
DailyOilGasPrices 188 30 2 Economy (25%/23%) (10%/8%)
DhakaHourlyAirQuality 2068 24 2 Environment (2%/3%) (0%/2%)
ElectricityPredictor 810 168 2 Energy (4%/5%) (7%/13%)
FloodModeling3 613 266 2 Environment (20%/22%) (20%/31%)
HouseholdPowerConsumption1 1431 1440 2 Energy (4%/7%) (14%/27%)
HotwaterPredictor 351 168 2 Energy (1%/1%) (4%/5%)
MadridPM10Quality 6923 168 2 Environment (4%/5%) (11%/16%)
ParkingBirmingham 1888 14 2 Environment (7%/28%) (5%/22%)
PrecipitationAndalusia 672 365 2 Environment (0%/1%) (3%/4%)
SierraNevadaMountainsSnow 500 30 2 Environment (6%/7%) (-2%/4%)
SolarRadiationAndalusia 672 365 2 Energy (1%/1%) (4%/4%)
SteamPredictor 300 168 2 Energy (2%/2%) (3%/6%)
TetuanEnergyConsumption 364 144 2 Energy (5%/5%) (1%/6%)
WindTurbinePower 852 144 2 Energy (11%/12%) (15%/21%)
Appendix F Supplementary results
F.1 Additional figures : Standard cost setting
F.2 Additional figures : Anomaly detection cost setting
F.3 Removing calibration
F.4 Ablation study on non-myopic trigger models
One of the key component when it comes to trigger functions is the non-myopic ability to anticipate the likely future. In this section, we conduct an ablation study over the anticipation-based property by making those type of approaches myopic. In particular, Economy and Calimera will be of interest here. More precisely, the expected cost horizon has been limit to 1 for Economy. Concerning Calimera, the backward cost propagation has been blocked such that the only available information to trigger be the expected cost difference between current and next timestamp.
F.5 Changing the base classifier
classifier
$\alpha$ MiniROCKET Weasel 2.0 tsfresh&XGBoost
0 12.60% 11.95% 16.36%
0.1 31.56% 18.44% 37.01%
0.2 32.86% 18.96% 36.49%
0.3 36.49% 19.48% 32.34%
0.4 39.87% 17.53% 31.30%
0.5 43.12% 19.10% 26.62%
0.6 42.60% 23.25% 23.51%
0.7 44.16% 24.94% 20.52%
0.8 48.44% 25.45% 15.71%
0.9 48.70% 26.62% 14.29%
1 42.21% 26.75% 14.68%
F.6 Impact of z-normalization
@@
Clearly, using z-normalized datasets is not applicable in practice, as it would require knowledge of the entire incoming time series. In a research context, previous work has used such training sets to test the proposed algorithms. Our goal here, is to assess whether this could have a large impact on the performances. For example, when a normalized time series has a low variance at the beginning, we can expect a high variance in the rest of the series since the mean variance is 1. There is therefore an information leakage that can be exploited by an ECTS algorithm, while this is not representative of what happens in real applications. A proposal such as the one presented by Schäfer & Leser (2020), where the z-normalization of available time series is repeated at each time step, has its own problems. In particular, it means that if a single classifier is used for all time steps, the representation of $\mathbf{x}_{t}$ can be different at times $t$ and $t+1$ and all further time steps which can induce confusion for the classifier.
@@
On the one hand, z-normalization induces an information leakage that could help methods to unduly exploit knowledge about the future of incoming time series. On the other hand, any normalization rescales the signal and therefore, potentially, hinder the recognition of telltale features. So, does z-normalization affect the performance of ECTS methods? And if yes, in which way?
@@
In order to answer this question, we took the new datasets collection described in Section 4.1. They are indeed not z-normalized originally. We duplicate and z-normalized them to get a second collection. As explained in Section 4.1, some extrinsic regression datasets have been converted into classification ones. Here, the threshold value chosen to discretize the output into binary classes has been set to the median of the regression target. In this way, classes within those datasets are equally populated.
@@
In these experiments, the delay cost is linear as in Section 4.2 and as in most of the literature. Figure LABEL:fig:heat_bar reports pairwise comparisons done on the 35 datasets. We look at $\alpha=0.8$, as this is the only value for which significant differences are observed. One can see that most of the trigger models do not actually benefit from the z-normalization. Quite the opposite: out of nine trigger models, only one, i.e. Ecdire, actually has a better mean AvgCost when being trained on z-normalized data. Regarding the remaining methods, both Stopping Rule and TeaserAvg perform significantly worse when operating on z-normalized data. Those trends are quite similar for other $\alpha$ values, without any significance on the statistical tests though. Thus, while z-normalization has some impact, since privileged information from the future can be leaked, our experiments, for the proposed datasets collection at least, show that this does not alter the overall results reported in the literature, and are globally in accordance with the results presented in Section 4.
F.7 Anomaly detection cost setting : an ablation study